The width of a rectangular room is 2 m shorter than its length. If the perimeter of the room is 72 m, calculate the length and width of the room
Answer by ikleyn(52781)

.

    W     + L = 72/2 = 36


    (L-2) + L = 36


     2L - 2 = 36

     2L     = 36 + 2 = 38

      L              = 38/2 = 19.

      W              = 19 - 2 = 17.


ANSWER.  The dimensions of the rectangle are  17 m and 19 m.

Solved.